OpenStudy (anonymous):

Math Analysis: Prove the given identity tan x/1 + sec x + 1 + sec x/ tan x = 2 csc x

OpenStudy (anonymous):

another guess \[\frac{\tan(x)}{1+\sec(x)}+\frac{1+\sec(x)}{\tan(x)}=2\csc(x)\] maybe?

OpenStudy (anonymous):

yeah!! :DD

OpenStudy (anonymous):

as with most of these i would get rid of the trig and do the algebra first

OpenStudy (anonymous):

make \[a=\cos(x), b=\sin(x)\] and see what you get with the "a" and "b"

OpenStudy (anonymous):

is it clear what i mean?

OpenStudy (anonymous):

no. lol

OpenStudy (anonymous):

tangent is sine over cosine, so where you see tan put \[\frac{b}{a}\] and secant is \[\frac{1}{\cos(x)}\] so replace it by \[\frac{1}{a}\] that will demystify the algebra a bit

OpenStudy (anonymous):

ok i did that..

OpenStudy (anonymous):

\[\frac{\frac{b}{a}}{1+\frac{1}{a}}+\frac{1+\frac{1}{a}}{\frac{b}{a}}=\frac{2}{b}\] is what you want. there may be a trig step involved but do the algebra first. on the left hand side multiply top and bottom by "a"

OpenStudy (anonymous):

you get \[\frac{b}{a+1}+\frac{a+1}{b}=\frac{2}{b}\] so we can add the fractions and we are almost done

OpenStudy (anonymous):

\[\frac{b^2+a^2+2a+1}{b(a+1)}=\frac{2}{b}\] and now comes the only trig step of the whole problem

OpenStudy (anonymous):

since \[\sin^2(x)+\cos^2(x)=1\] i can replace \[a^2+b^2\] by 1 and get \[\frac{2a+2}{b(a+1)}=\frac{2}{b}\] \[\frac{2(a+1)}{b(a+1)}=\frac{2}{b}\] \[\frac{2}{b}=\frac{2}{b}\]

OpenStudy (anonymous):

if you like you can rewrite all of this replacing "a" by cosine and "b" by sine and you get the same thing

OpenStudy (anonymous):

ty your answers were very thorough :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

btw wouldnt cos x or (a) cancel at the bottom of (b/a)/1+1/a since we're multiplying both top and bottom by (a)???

OpenStudy (anonymous):

i didnt get how you got sin x/cos x +1

OpenStudy (anonymous):

how do i get sin x/cos x +1 doesn't cos x on the bottom cancel out?

OpenStudy (bahrom7893):

It's plus, not multiply.

OpenStudy (anonymous):

no. he said multiply top and bottom by a!!!

OpenStudy (anonymous):

a = cos

OpenStudy (anonymous):

when i multiply cos x to the bottom (1 + 1/cosx) shouldn't cosx be canceled out?

OpenStudy (anonymous):

i dont get how he got sinx/cos x +1

OpenStudy (bahrom7893):

Are u talkin about this step?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what step is confusing?

OpenStudy (anonymous):

wouldn't cos x cancel out on the botom of sinx/cosx/1+1/cosx?

OpenStudy (anonymous):

you mean here \[\frac{\frac{b}{a}}{1+\frac{1}{a}}+\frac{1+\frac{1}{a}}{\frac{b}{a}}=\frac{2}{b}\]

OpenStudy (anonymous):

so shouldn't it have been sin x/2??? instead of sinx/cosx +1

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

when you multiply top and bottom by "a" you get \[\frac{b}{a+1}+\frac{a+1}{b}=\frac{2}{b}\]

OpenStudy (anonymous):

i was waiting for you to finish explaining your answer to that Mikey guy, im sorry but he's incredibly stupid I can see why you raged quit on him

OpenStudy (anonymous):

yes but the a would cancel out when you multiply the bottom by a

OpenStudy (anonymous):

is that step clear or no?

OpenStudy (anonymous):

let me write it step by step

OpenStudy (anonymous):

NOO!! I feel like your not explaining enough D:

OpenStudy (rogue):

Lilfayfay, stop being rude & disrespectful, don't go questioning people's intelligence and calling people "incredibly stupid." You couldn't even put the parenthesis in the correct places and your the one saying I'm not very good at math...

OpenStudy (anonymous):

\[\frac{\frac{b}{a}}{1+\frac{1}{a}}=\frac{\frac{b}{a}}{1+\frac{1}{a}}\times \frac{a}{a}\] \[=\frac{a(\frac{b}{a})}{a(1+\frac{1}{a})}\] \[=\frac{b}{a+1}\]

OpenStudy (anonymous):

could you please hurry explaining i need to go do Ap Chem h.w

OpenStudy (anonymous):

you have to multiply everything in the denominator by "a" that is why you get \[\frac{b}{a+1}\]

OpenStudy (anonymous):

@rogue im sorry but im just taking really hard classes this year, im getting really frustrated with Math atm

OpenStudy (rogue):

I took the same classes last year, deal with it.

OpenStudy (anonymous):

so the "a" in the numerator cancels, and the "a" in the denominator cancels as well, but you have to multiply the "1" by "a" also

OpenStudy (anonymous):

i bet you did.. and i wasn't being rude ur just not explaining well thats all! :\

OpenStudy (anonymous):

ok ty <333 You're the best!!! :DDD

OpenStudy (rogue):

How could I explain if I didn't know what the correct question was? Nevermind, good luck with your homework. It's a lot of work, but try not to get frustrated...

OpenStudy (anonymous):

i already showed you what the question was, I even put it in parenthesis!!! look back I put it in parenthesis for you :P

OpenStudy (rogue):

"(tan x/1) + sec x + 1 + (sec x/ tan x) = 2 csc x" This is not the same as what Sat had. Sat had (tan x)/(1 + sec x) + (1 + sec x)/(tan x) =P

OpenStudy (anonymous):

yes it is lol you just cant read equations properly like he can (no offense) :p

OpenStudy (rogue):

No point in arguing with a girl; they always win. Good fight. You win.